Two geological field teams are working in a remote area. A global positioning system (GPS) tracker at their base camp shows the location of the first team as 39 km away, 22° north of west, and the second team as 30 km away, 38° east of north. When the first team uses its GPS to check the position of the second team, what does the GPS give for the following?

(a) the second team's distance from the first team

(b) the second team's direction from the first team, measured relative to due east
Answer #1

deg 22 deg

Let \vec{A} represent location of first team from base

Ā= 39km at 22° north of west

A = 39km at 158° from east

Ā= Ari + Ayj = Acosti + Asin j = 39cos158i + 39 sin 158j

A = -36.161 +14.617

====================

Let \vec{B} represent location of second team from first team

B = Bi + Bj

=============

Let \vec{R} represent location of second team from base

R= 30km at 38° east of north

\vec{R}=30km\ at\ 52^{\circ}\ from\ east

\vec{R}=R_{x}i+R_{y}j=Rcos\theta i+Rsin\theta j=30cos52i+30sin52j

{\color{Red} \vec{R}=18.47i+23.64j}

====================

R = A+B

(18.47i+23.64j)=(-36.16i+14.61j)+(B_{x}i+B_{y}j)

18.47i+23.64= -36.161 +14.61; + Bi + Buj

18.47i+23.64j=(-36.16+B_{x})i+(14.61+B_{y})j

Equate i terms

18.47=-36.16+B_{x}

18.47+36.16=B_{x}

{\color{Red} B_{x}=54.63}

============

Equate j terms

23.64=14.61+B_{y}

23.64-14.61=B_{y}

{\color{Red} B_{y}=9.03}

=================

(a) Magnitude

B=\sqrt{B_{x}^{2}+B_{y}^{2}}=\sqrt{54.63^{2}+9.03^{2}}

ANSWER :{\color{Red} B=55.37km}

======================

(b) Direction

( ) 2014 = 0

\theta =tan^{-1}(\frac{9.03}{54.63})

ANSWER :{\color{Red} \theta =9.39^{\circ}\ from\ east}

===========================